# Cointegration test in time series analysis by global optimisation

### v2

@@ 4.2. Test for higher dimensional finite samples
(4.3) $\begin{pmatrix}s_{1t}\\ s_{2t}\\ s_{3t}\end{pmatrix}-\begin{pmatrix}-0.3&-0.03&-0.375\\ -1.88&-1.472&-2.08\\ 1.07&0.428&1.27\end{pmatrix}\begin{pmatrix}s_{1,t-1}\\ s_{2,t-1}\\ s_{3,t-1}\end{pmatrix}=\begin{pmatrix}\varepsilon_{1t}\\ \varepsilon_{2t}\\ \varepsilon_{3t}\end{pmatrix}-\begin{pmatrix}0.75&0.075&0.9375\\ 4.7&3.68&5.2\\ -2.675&-1.07&-3.175\end{pmatrix}\begin{pmatrix}\varepsilon_{1,t-1}\\ \varepsilon_{2,t-1}\\ \varepsilon_{3,t-1}\end{pmatrix}.$
@@
(4.4) $\begin{pmatrix}z_{1t}\\ z_{2t}\\ z_{3t}\end{pmatrix}-\begin{pmatrix}13&12&14\\ 10&1&12.5\\ 0&0&0\end{pmatrix}\begin{pmatrix}z_{1,t-1}\\ z_{2,t-1}\\ z_{3,t-1}\end{pmatrix}=\begin{pmatrix}\delta_{1t}\\ \delta_{2t}\\ \delta_{3t}\end{pmatrix}-\begin{pmatrix}-7.3125&-6.75&-7.875\\ -5.625&-0.5625&-7.03125\\ 0&0&0\end{pmatrix}\begin{pmatrix}\delta_{1,t-1}\\ \delta_{2,t-1}\\ \delta_{3,t-1}\end{pmatrix}.$
@@
$\hat{W}=\begin{pmatrix}-0.0402&0.9856&-0.1640\\ 0.4652&-0.1268&-0.8761\\ 0.8843&0.1115&0.4534\end{pmatrix}.$
@@ 5. Conclusion
Furthermore, we have introduced how the maximization of nongaussianity can be applied to test cointegration, which goes beyond bivariate cointegration. This method has shown its effectiveness in discovering the underlying stationary components of a mixed observed time series, thereby facilitating the subsequent process of finding cointegration. Generally, its preponderance clearly lies in its applicability to higher dimensional scenarios and overall satisfactory and promising accuracy.
@@ 7. Appendix
$\text{H}_{0}:\mbox{rank}\left(\Pi\right)=r_{0}\text{ against }\text{H}_{1}:\mbox{rank}\left(\Pi\right)=r_{0}+1$

### v3

@@ 4.1. Test for unit-root nonstationary simulated series
In this section, we carry out a Monte Carlo experiment with 1000 repetitions on a simulated time series as a demonstration of our methods. We will also use bias and mean square error (MSE) for comparison in different sample sizes.
@@ 4.2. Monte Carlo Simulation Results
A Monte Carlo simulation is then carried out to compare the performance of our methods and Johansen’s in different sample sizes. As shown in Table 3.
Table 3. Monte Carlo Simulation Results: Bias and MSE for Cointegration Vector Estimation
Sample Size (T) Method Bias($\beta_{1}$) MSE($\beta_{1}$)
10 Decorrelation $6.3220$ $4.64\text{\times}{10}^{03}$
Max. Nongaussianity $1.4708$ $4.05\text{\times}{10}^{01}$
Johansen $2.2996$ $1.00\text{\times}{10}^{02}$
15 Decorrelation $4.0491$ $1.85\text{\times}{10}^{03}$
Max. Nongaussianity $0.9306$ $3.00\text{\times}{10}^{00}$
Johansen $1.3117$ $4.84\text{\times}{10}^{01}$
20 Decorrelation $2.0884$ $5.79\text{\times}{10}^{01}$
Max. Nongaussianity $1.0794$ $1.04\text{\times}{10}^{02}$
Johansen $1.2554$ $2.48\text{\times}{10}^{02}$
25 Decorrelation $3.6321$ $2.98\text{\times}{10}^{03}$
Max. Nongaussianity $0.7921$ $3.62\text{\times}{10}^{01}$
Johansen $0.3454$ $2.08\text{\times}{10}^{00}$
30 Decorrelation $1.5528$ $4.47\text{\times}{10}^{01}$
Max. Nongaussianity $0.5343$ $1.37\text{\times}{10}^{00}$
Johansen $0.2072$ $3.66\text{\times}{10}^{-01}$
40 Decorrelation $0.7040$ $2.61\text{\times}{10}^{01}$
Max. Nongaussianity $0.3777$ $2.66\text{\times}{10}^{-01}$
Johansen $0.1498$ $9.60\text{\times}{10}^{-01}$
50 Decorrelation $0.5428$ $3.05\text{\times}{10}^{01}$
Max. Nongaussianity $0.3048$ $1.82\text{\times}{10}^{-01}$
Johansen $0.0906$ $4.32\text{\times}{10}^{-02}$
100 Decorrelation $0.1355$ $7.10\text{\times}{10}^{-01}$
Max. Nongaussianity $0.1809$ $8.47\text{\times}{10}^{-02}$
Johansen $0.0367$ $2.62\text{\times}{10}^{-03}$
3000 Decorrelation $0.0022$ $1.34\text{\times}{10}^{-05}$
Max. Nongaussianity $0.0089$ $3.85\text{\times}{10}^{-04}$
Johansen $0.0011$ $2.18\text{\times}{10}^{-06}$
Figure 3. Monte Carlo Simulation Error Comparison
@@ 4.3. Test for higher dimensional finite samples
(4.3) $\begin{pmatrix}s_{1t}\\ s_{2t}\\ s_{3t}\end{pmatrix}-\begin{pmatrix}-0.3&-0.03&-0.375\\ -1.88&-1.472&-2.08\\ 1.07&0.428&1.27\end{pmatrix}\begin{pmatrix}s_{1,t-1}\\ s_{2,t-1}\\ s_{3,t-1}\end{pmatrix}=\begin{pmatrix}\varepsilon_{1t}\\ \varepsilon_{2t}\\ \varepsilon_{3t}\end{pmatrix}-\begin{pmatrix}0.75&0.075&0.9375\\ 4.7&3.68&5.2\\ -2.675&-1.07&-3.175\end{pmatrix}\begin{pmatrix}\varepsilon_{1,t-1}\\ \varepsilon_{2,t-1}\\ \varepsilon_{3,t-1}\end{pmatrix}$
@@
(4.4) $\begin{pmatrix}z_{1t}\\ z_{2t}\\ z_{3t}\end{pmatrix}-\begin{pmatrix}13&12&14\\ 10&1&12.5\\ 0&0&0\end{pmatrix}\begin{pmatrix}z_{1,t-1}\\ z_{2,t-1}\\ z_{3,t-1}\end{pmatrix}=\begin{pmatrix}\delta_{1t}\\ \delta_{2t}\\ \delta_{3t}\end{pmatrix}-\begin{pmatrix}-7.3125&-6.75&-7.875\\ -5.625&-0.5625&-7.03125\\ 0&0&0\end{pmatrix}\begin{pmatrix}\delta_{1,t-1}\\ \delta_{2,t-1}\\ \delta_{3,t-1}\end{pmatrix}$
@@
$\hat{W}=\begin{pmatrix}-0.0402&0.9856&-0.1640\\ 0.4652&-0.1268&-0.8761\\ 0.8843&0.1115&0.4534\end{pmatrix}$
@@
4.4. High-Dimensional Example with Multiple Cointegration Vectors
To demonstrate the capability of our method in higher-dimensional settings with multiple cointegration relationships, we consider a 4-dimensional system with two cointegration vectors. The data generating process is defined as follows:
Let $\mathbf{c}_{t}=(c_{1t},c_{2t},c_{3t},c_{4t})^{\top}$ where:
$c_{1t}$ $=c_{1,t-1}+\varepsilon_{1t},\quad\varepsilon_{1t}\sim\mathscr{N}(0,0.01)$
$c_{2t}$ $=c_{2,t-1}+\varepsilon_{2t},\quad\varepsilon_{2t}\sim\mathscr{N}(0,0.01)$
$c_{3t}$ $\sim t(5)$
$c_{4t}$ $\sim t(5)$
The components $\mathbf{c}_{t}$ are linearly mixed through a random invertible matrix $A\in\mathbb{R}^{4\times 4}$ to produce the observed series:
$\mathbf{s}_{t}=A\mathbf{c}_{t}$
We generate $T=3000$ observations after a burn-in period of 500 points. The true cointegration vectors correspond to the last two rows of $A^{-1}$, which represent the stationary non-Gaussian components ($c_{3t},c_{4t}$).
Applying the maximization of nongaussianity method (Algorithm 2, Section 3.3) with $F(x)=\ln\cosh(x)$, we obtain the demixing matrix $\hat{W}$. The recovered components are tested for stationarity using the Augmented Dickey-Fuller test at 1% significance level. The two components with the most negative test statistics (strongest evidence of stationarity) are identified as cointegration relationships.
Again, standard critical values may lead to size distortion due to applying the ADF on time series resulting from the optimisation procedure. Hence, a more negative value should be used for residuals than usual. In our particular case, the critical value for both cointegration vectors are -11.3892 and -10.3729 respectively, which are widely more negative than usual critical values, while a rigorous analysis on critical values used for BSS procedures should be implemented via a wider range of simulation.
Results from a representative simulation show:
Estimated cointegration vectors:
$\hat{\bm{\beta}}_{1}=[1.0000,0.6356,0.4620,-0.2958]^{\top}\quad$
$\hat{\bm{\beta}}_{2}=[1.0000,0.2799,0.8265,0.1067]^{\top}\quad$
True cointegration vectors:
$\bm{\beta}_{1}=[1.0000,0.6339,0.4728,-0.2852]^{\top}$
$\bm{\beta}_{2}=[1.0000,0.3021,0.8325,0.1062]^{\top}$
The relative errors in the coefficients are below $0.05\%$, confirming the method accurately recovers multiple cointegration vectors in high dimensions. Figure 6 visualizes the components and their autocorrelation functions.
Figure 4. Recovered components via maximization of nongaussianity. Components 3 and 4 (bottom) show stationarity and match the true non-Gaussian sources.
@@ 5. Conclusion
Furthermore, we have introduced how the maximisation of nongaussianity can be applied to test cointegration, which goes beyond bivariate cointegration. This method has shown its effectiveness in discovering the underlying stationary components of a mixed observed time series, thereby facilitating the subsequent process of finding cointegration. Generally, its preponderance clearly lies in its applicability to higher dimensional scenarios and overall satisfactory and promising accuracy.
@@ 7. Appendix
[…]